# Speculative End-Turn Detector for Efficient Speech Chatbot Assistant

### v1

@@ Abstract
Spoken dialogue systems powered by large language models have demonstrated remarkable abilities in understanding human speech and generating appropriate spoken responses. However, these systems struggle with end-turn detection (ETD)—the ability to distinguish between user turn completion and hesitation. This limitation often leads to premature or delayed responses, disrupting the flow of spoken conversations. In this paper, we introduce the ETD Dataset, the first public dataset for end-turn detection. The ETD dataset consists of both synthetic speech data generated with text-to-speech models and real-world speech data collected from web sources. We also propose SpeculativeETD, a novel collaborative inference framework that balances efficiency and accuracy to improve real-time ETD in resource-constrained environments. Our approach jointly employs a lightweight GRU-based model, which rapidly detects the non-speaking units in real-time on local devices, and a high-performance Wav2vec-based model running on the server to make a more challenging classification of distinguishing turn ends from mere pauses. Experiments demonstrate that the proposed SpeculativeETD significantly improves ETD accuracy while keeping the required computations low. Datasets and code will be available after the review.
@@ 1 Introduction
Recent advancements in large language models (LLMs) have spurred extensive research into various LLM-based agents (Song et al., 2023; Shao et al., 2024; Zhang et al., 2024; Xie et al., 2024). Many studies focus on integrating LLMs with spoken dialogue systems to enable more engaging human-computer interaction (Mitsui et al., 2023; Yan et al., 2024; Ma et al., 2024; Veluri et al., 2024). However, LLM-based spoken dialogue systems face a critical challenge: when the user suddenly stops speaking, the systems often fail to discern whether the user has completed their turn or is merely pausing to think during a conversation (Lin et al., 2022; Umair et al., 2024). For instance, as illustrated in Figure 1, when a user momentarily pauses while contemplating a specific question, the LLM may erroneously interpret this pause as the end of the turn, leading to premature and off-target responses. This task, called end-turn detection (ETD), is a challenging task that requires nuanced mechanisms based on a firm understanding of the content of the speech; naïve cues, such as the pause duration, fails to provide meaningful information to distinguish between the turn completion and user hesitation (Ten Bosch et al., 2005).
@@
To tackle this bottleneck, we construct and release the ETD dataset, an open-to-public dataset that is specifically designed for training and evaluating end-turn detection models. The ETD dataset consists of more than 120k samples with over 300 hours of conversational data. In particular, the dataset consists of both synthetic spoken conversation generated via applying text-to-speech (TTS) generation on text dialogue dataset and real-world dialogues collected from web sources.
@@
| GRU | M | 79.34 |
| Wav2vec 2.0 | M | 99.49 |
Beyond dataset construction, we propose an efficient end-turn detection algorithm tailored for real-time chatbot interactions under resource-constrained environments. While transformer-based audio processing models achieve strong performance, their high computational cost makes them impractical to be used in highly repetitive tasks with real-time applications, such as ETD. In contrast, models under 1M parameters can run efficiently on local devices (Qi et al., 2023; Wu & Chen, 2023) but suffer from degraded accuracy (see Table 1). To address this gap, we introduce SpeculativeETD, a novel collaborative inference framework that balances efficiency and accuracy. In this framework approach, we adopt both the lightweight GRU-based model (Chung et al., 2014) and the high-performance Wav2vec-based model (Baevski et al., 2020). Here, the lightweight model operates on local devices in real time and keeps track of whether the speaker is speaking or not (either as a pause or end-of-turn). Whenever the model detects the silence, the model queries the high-performance model to make a more challenging prediction, i.e., telling whether the silence marks the end of the speaker’s turn or not. This framework removes the need for the high-performance model to be running in real-time, dramatically saving the amount of required computations with minimal degradation in accuracy.
@@ 2 Related work
Turn-taking plays a crucial role in human communication, allowing speakers to transition smoothly without interruptions. Its detection is guided by linguistic, prosodic, and non-verbal cues such as pitch variation, speech rhythm, and gaze direction (Gravano & Hirschberg, 2011; Levinson & Torreira, 2015). Early turn-taking models have used finite-state machines to predict turn timing and duration (Raux & Eskenazi, 2009).
In spoken dialogue systems, turn-taking ensures fluid human-computer interactions. Various approaches have been explored for turn boundary prediction, including end-to-end models based on automatic speech recognition (ASR) (Chang et al., 2022) and transformer-based models incorporating contextual and pragmatic cues (Ekstedt & Skantze, 2020). More recently, full-duplex LLM-based agents have been developed to enable simultaneous listening and responding (Défossez et al., 2024). However, existing models remain computationally expensive and rely on private datasets, limiting real-time applicability and reproducibility. Our approach addresses these challenges by introducing an efficient turn-detection method and an open-source dataset.
@@
Disfluencies, such as interruptions, filler words, or self-corrections, occur frequently in spontaneous speech and can disrupt the natural flow of conversations. Identifying and handling these phenomena is crucial for maintaining coherent interactions in spoken dialogue systems. Traditional approaches primarily rely on text-based detection, analyzing transcriptions with span classification models to improve accuracy (Ghosh et al., 2022). Although effective, these methods depend on ASR, which introduces latency and errors. Recent studies have explored direct speech-based disfluency detection (Zhou et al., 2024) and integrated frameworks that combine ASR with disfluency removal for smoother real-time processing (Lou & Johnson, 2020). However, these methods often struggle with spontaneous speech variations and real-time efficiency.
@@
3 End-Turn Detection (ETD) Dataset
@@
The ETD dataset consists of multiple audio speech files labeled with the ternary segmentation information. That is, the label of each audio file is a sequence of the (state, starting time) pair of each same-state segment, e.g.,
We construct the ETD dataset using both synthetic (for both training and evaluation) and real (for evaluation only) conversational audio segments.
3.1 Synthetic data generation
We have constructed the synthetic audio dataset of multi-turn conversation by applying the TTS (text-to-speech) model to the existing text dialogue dataset. In particular, we have used the MultiWOZ corpus (Budzianowski et al., 2018) as the base dataset. As the TTS model, we have used multiple TTS models, in particular the MeloTTS (Zhao et al., 2023) and Google Cloud Text-to-Speech, to ensure the diversity of the generated speech.
An important thing to note here is that a naïve text-to-speech translation of text dialogues will generate data consisting only of speaking units and gaps, as the textual cues for the “pause” are typically missing in a text dialogue. Thus, we artificially insert the pauses to the conversation in three different styles (Figure 2(a)):
Base: We do not insert any pause.
w/ Pause: We randomly select tiny hesitation—typically inserted by TTS models to mimic human breathe (Braunschweiler & Buchholz, 2011; Hwang et al., 2023; Yang et al., 2024)—and extend it into a pause. The duration of the pause has been randomly drawn from the distribution (where the unit is in seconds). For some randomly picked samples, we have also removed the segments that follow the pause, generating samples without a gap at the end of the audio file.
3.2 Real data generation
We have also collected real spoken conversation data from YouTube and Buckeye speech corpus (Pitt et al., 2005); we have selected these sources as they contain many spontaneous, natural dialogues between two speakers with frequent turn-takings. To ensure that the collected samples involve exactly two speakers, we have applied a pretrained speaker diarization toolkit (Bredin, 2023) to each audio file and filtered out those with more than two identified speakers.
Given conversation data, we again utilized the speaker diarization tool to generate segmentations of the speech of two different speakers (Figure 2 (b)). Any silent interval between speech segments that exceed 200ms is labeled as either pause or gap, depending on whether the neighboring speech segments belong to the same speaker or not; here, the 200ms criterion follows the commonly used thresholds in the turn-taking literature (Stivers et al., 2009; Heldner & Edlund, 2010; Nguyen et al., 2022; Li et al., 2022). Similarly to the synthetic dataset, we have randomly selected some samples and removed the audio segment that follows a randomly chosen gap or pause.
3.3 Other details
Splits and statistics. The default train/dev/test split of the synthetic data follows that of the MultiWOZ dataset (Budzianowski et al., 2018). The statistics of the synthetic and real datasets are given in Table 8 and Table 9, respectively.
@@ 3.3 Other details
| # Total duration (h) | 200.86 | 40.91 | 89.74 | 70.21 |
| # Average duration (s) | 5.90 | 5.19 | 6.87 | 5.37 |
@@
| # Total duration (h) | 158.4 | 32.14 | 70.8 | 55.46 |
| # Average duration (s) | 5.89 | 5.18 | 6.84 | 5.37 |
@@
| # Total duration (h) | 21.25 | 4.34 | 9.48 | 7.43 |
| # Average duration (s) | 5.96 | 5.20 | 6.94 | 5.44 |
@@
| # Total duration (h) | 21.21 | 4.43 | 9.46 | 7.32 |
| # Average duration (s) | 5.93 | 5.24 | 6.93 | 5.36 |
@@
| # Samples | 1,665 | 425 | 190 | 1,050 | - | - |
| Duration (h) | 32.25 | 3.26 | 0.36 | 28.68 | 6.33 | 3.24 |
@@ 4 Method
For a smooth real-time conversation, one needs to run the end-turn detector repetitively at a rapid frequency. Thus, it is desirable to deploy an ETD model with an extremely small computational footprint. However, lightweight models—e.g., a recurrent model with a single GRU layer—tend to suffer from a low detection accuracy. On the other hand, heavier transformer-based models typically require a large computational cost despite being highly accurate in the end-turn detection (see Section 6).
To enjoy the best of both worlds, the SpeculativeETD is designed as a collaborative inference framework that utilizes both a lightweight model (say, GRU), which runs on-device, and a heavier model (say, Wav2vec 2.0), which runs on the server. In a nutshell, the proposed SpeculativeETD operates in two stages (Figure 3):
@@
The key advantages of this framework are twofold: (1) The trickier decision that requires a large computation (i.e., telling “Gap” vs. “Pause”) can happen only once per each consecutive segment of silence, not at every time frame. As we will see in Section 6, this helps save the computation more than 10 with only a small degradation prediction accuracy. (2) The two-stage pipeline allows us to utilize on-device computation by placing the lightweight model on the edge. By doing so, we no longer need the communication between the edge device and the server to be done in a continuous manner; instead, we can communicate once every silence.
@@
During real-time operation, the on-device model processes each new audio frame in constant time. Predictions are queued locally, and when predicted as ”not speaking,” they are sent to the server. The server-side model then processes the segment, determining whether an end-turn event has occurred or if streaming should continue.
@@ 5 Experimental Setup
We now empirically validate the usefulness of the constructed ETD dataset (Section 3) and the effectiveness of the proposed method, SpeculativeETD (Section 4).
Task. We consider two scenarios of the ETD task:
@@
VAP (Ekstedt & Skantze, 2022): A popular open-source pretrained turn-taking model, which can also be employed for end-turn detection tasks. We use the pretrained encoder as frozen and train the predictor module of the model from scratch using the ETD dataset. We have used the “comparative” predictor head, which fits the ETD task.
GRU (Chung et al., 2014): We construct a lightweight GRU model with 1M parameters, with two Conv2D layers and one GRU layer, and train from scratch with the ETD dataset.
Wav2vec 2.0 (Baevski et al., 2020): A relatively large transformer-based model with 94M parameters, which achieves high performance in many speech processing tasks. We fine-tune the full model using the ETD dataset.
@@
In all experiments, we have trained the models using the training split of the synthetic data in the ETD dataset. For evaluating the models, we have used both the test split of the synthetic dataset and the full split of the real dataset.
@@
We have used the AdamW optimizer (Loshchilov. & Hutter, 2019) for all tasks and models and trained for 10 epochs. The learning rate has been selected with a random search over the interval , using the validation split of the synthetic ETD dataset. The weight decay values are also randomly searched in the range . The batch size has been tuned over for GRU, for VAP, and for Wav2vec 2.0. The searched hyperparameters are provided in Appendix C.

### v2

@@ Abstract
Spoken dialogue systems powered by large language models have demonstrated remarkable abilities in understanding human speech and generating appropriate spoken responses. However, these systems struggle with end-turn detection (ETD)—the ability to distinguish between user turn completion and hesitation. This limitation often leads to premature or delayed responses, disrupting the flow of spoken conversations. In this paper, we introduce the OpenETD Dataset, the first public dataset for end-turn detection. The OpenETD dataset consists of both synthetic speech data generated with text-to-speech models and real-world speech data collected from web sources. We also propose SpeculativeETD, a novel collaborative inference framework that balances efficiency and accuracy to improve real-time ETD in resource-constrained environments. Our approach jointly employs a lightweight GRU-based model, which rapidly detects the non-speaking units in real-time on local devices, and a high-performance Wav2vec-based model running on the server to make a more challenging classification of distinguishing turn ends from mere pauses. Experiments demonstrate that the proposed SpeculativeETD significantly improves ETD accuracy while keeping the required computations low.
Speculative End-Turn Detector for Efficient Speech Chatbot Assistant
Hyunjong Ok1,2 Suho Yoo2,3 Jaeho Lee1 1POSTECH 2HJ AILAB 3KAIST {hyunjong.ok, uso7d0}@gmail.com, jaeho.lee@postech.ac.kr
@@ 1 Introduction
Recent advancements in large language models (LLMs) have spurred extensive research into various LLM-based agents (Song et al., 2023; Shao et al., 2024; Zhang et al., 2024; Xie et al., 2024). Many studies focus on integrating LLMs with spoken dialogue systems to enable more engaging human-computer interaction (Mitsui et al., 2023; Yan et al., 2024; Ma et al., 2024; Veluri et al., 2024). However, LLM-based spoken dialogue systems face a critical challenge: when the user suddenly stops speaking, the systems often fail to discern whether the user has completed their turn or is merely pausing to think during a conversation (Lin et al., 2022; Umair et al., 2024). For instance, as illustrated in Figure˜1, when a user momentarily pauses while contemplating a specific question, the LLM may erroneously interpret this pause as the end of the turn, leading to premature and off-target responses. This task, called end-turn detection (ETD), is a challenging task that requires nuanced mechanisms based on a firm understanding of the content of the speech; naïve cues, such as the pause duration, fails to provide meaningful information to distinguish between the turn completion and user hesitation (Ten Bosch et al., 2005).
@@
To tackle this bottleneck, we construct and release the OpenETD dataset, an open-to-public dataset that is specifically designed for training and evaluating end-turn detection models. The OpenETD dataset consists of more than 120k samples with over 300 hours of conversational data. In particular, the dataset consists of both synthetic spoken conversation generated via applying text-to-speech (TTS) generation on text dialogue dataset and real-world dialogues collected from web sources.
@@
| GRU | M | 79.7 |
| Wav2vec 2.0 | M | 99.3 |
Beyond dataset construction, we propose an efficient end-turn detection algorithm tailored for real-time chatbot interactions under resource-constrained environments. While transformer-based audio processing models achieve strong performance, their high computational cost makes them impractical to be used in highly repetitive tasks with real-time applications, such as ETD. In contrast, models under 1M parameters can run efficiently on local devices (Qi et al., 2023; Wu and Chen, 2023) but suffer from degraded accuracy (see Table˜1). To address this gap, we introduce SpeculativeETD, a novel collaborative inference framework that balances efficiency and accuracy. In this framework approach, we adopt both the lightweight GRU-based model (Chung et al., 2014) and the high-performance Wav2vec-based model (Baevski et al., 2020). Here, the lightweight model operates on local devices in real time and keeps track of whether the speaker is speaking or not (either as a pause or end-of-turn). Whenever the model detects the silence, the model queries the high-performance model to make a more challenging prediction, i.e., telling whether the silence marks the end of the speaker’s turn or not. This framework removes the need for the high-performance model to be running in real-time, dramatically saving the amount of required computations with minimal degradation in accuracy.
@@ 2 Related work
Turn-taking plays a crucial role in human communication, allowing speakers to transition smoothly without interruptions. Its detection is guided by linguistic, prosodic, and non-verbal cues such as pitch variation, speech rhythm, and gaze direction (Gravano and Hirschberg, 2011; Levinson and Torreira, 2015). Early turn-taking models have used finite-state machines to predict turn timing and duration (Raux and Eskenazi, 2009).
In spoken dialogue systems, turn-taking ensures fluid human-computer interactions. Various approaches have been explored for turn boundary prediction, including end-to-end models based on automatic speech recognition (ASR) (Chang et al., 2022) and transformer-based models incorporating contextual and pragmatic cues (Ekstedt and Skantze, 2020). More recently, full-duplex LLM-based agents have been developed to enable simultaneous listening and responding (Défossez et al., 2024). However, existing models remain computationally expensive and rely on private datasets, limiting real-time applicability and reproducibility. Our approach addresses these challenges by introducing an efficient turn-detection method and an open-source dataset.
@@
Disfluencies, such as interruptions, filler words, or self-corrections, occur frequently in spontaneous speech and can disrupt the natural flow of conversations. Identifying and handling these phenomena is crucial for maintaining coherent interactions in spoken dialogue systems. Traditional approaches primarily rely on text-based detection, analyzing transcriptions with span classification models to improve accuracy (Ghosh et al., 2022). Although effective, these methods depend on ASR, which introduces latency and errors. Recent studies have explored direct speech-based disfluency detection (Zhou et al., 2024) and integrated frameworks that combine ASR with disfluency removal for smoother real-time processing (Lou and Johnson, 2020). However, these methods often struggle with spontaneous speech variations and real-time efficiency.
@@
3 OpenETD: End-Turn Detection Dataset
@@
The OpenETD dataset consists of multiple audio speech files labeled with the ternary segmentation information. That is, the label of each audio file is a sequence of the (state, starting time) pair of each same-state segment, e.g.,
We construct the OpenETD dataset using both synthetic (for both training and evaluation) and real (for evaluation only) conversational audio segments.
@@
| Total duration (h) | 148.26 | 28.66 | 69.60 | 49.99 |
| Average duration (s) | 4.36 | 3.64 | 5.32 | 3.83 |
@@
| Total duration (h) | 116.83 | 22.49 | 54.87 | 39.47 |
| Average duration (s) | 4.35 | 3.63 | 5.30 | 3.82 |
@@
| Total duration (h) | 15.75 | 3.06 | 7.37 | 5.32 |
| Average duration (s) | 4.42 | 3.66 | 5.40 | 3.90 |
@@
| Total duration (h) | 15.68 | 3.11 | 7.36 | 5.20 |
| Average duration (s) | 4.39 | 3.69 | 5.39 | 3.81 |
@@
| # Samples | 2,898 | 1,113 | 135 | 1,650 | - | - |
| Duration (h) | 32.21 | 8.40 | 1.13 | 22.72 | 2.96 | 1.27 |
| Train | ||||||
| # Samples | 6,290 | 2,624 | 449 | 3,217 | - | - |
| Duration (h) | 117.16 | 31.12 | 5.47 | 80.58 | 5.54 | 2.46 |
| Dev | ||||||
| # Samples | 899 | 363 | 67 | 469 | - | - |
| Duration (h) | 16.23 | 3.60 | 0.91 | 11.72 | 0.74 | 0.34 |
| Test | ||||||
| # Samples | 1,798 | 769 | 133 | 896 | - | - |
| Duration (h) | 32.36 | 9.26 | 1.70 | 21.40 | 1.50 | 0.67 |
3.1 Synthetic data generation
We have constructed the synthetic audio dataset of multi-turn conversation by applying the TTS (text-to-speech) model to the existing text dialogue dataset. In particular, we have used the MultiWOZ corpus (Budzianowski et al., 2018) as the base dataset. As the TTS model, we have used Google Cloud Text-to-Speech.
An important thing to note here is that a naïve text-to-speech translation of text dialogues will generate data consisting only of speaking units and gaps, as the textual cues for the “pause” are typically missing in a text dialogue. Thus, we artificially insert the pauses to the conversation in three different styles (Figure˜2(a)):
Base: We do not insert any pause.
w/ Pause: We randomly select tiny hesitations—typically inserted by TTS models to mimic human breathing (Braunschweiler and Buchholz, 2011; Hwang et al., 2023; Yang et al., 2024)—and extend them into pauses. Here, we sample each pause duration from an Erlang distribution with shape and rate , truncated to . This yields a right-skewed distribution with mean and peak around –, matching the typical pattern of conversational pauses (many short, few long). Statistical validation against real data is provided in Appendix˜A. For some randomly picked samples, we also remove the segments that follow the pause, generating samples without a gap at the end of the audio file.
3.2 Real data generation
We have also collected real spoken conversation data from YouTube and Buckeye speech corpus (Pitt et al., 2005); we have selected these sources as they contain many spontaneous, natural dialogues between two speakers with frequent turn-takings. For YouTube, we used broad queries (e.g., “english conversation podcast”, “podcast”) via yt-dlp, retrieving up to relevance-ranked results per query, and kept only two-speaker conversations after diarization (monologues and multi-party discussions were filtered out). To ensure that the collected samples involve exactly two speakers, we have applied a pretrained speaker diarization toolkit (Bredin, 2023) to each audio file and filtered out those with more than two identified speakers.
Given conversation data, we again utilized the speaker diarization tool to generate segmentations of the speech of two different speakers (Figure˜2 (b)). Any silent interval between speech segments that exceed 200ms is labeled as either pause or gap, depending on whether the neighboring speech segments belong to the same speaker or not; here, the 200ms criterion follows the commonly used thresholds in the turn-taking literature (Stivers et al., 2009; Heldner and Edlund, 2010; Nguyen et al., 2022; Li et al., 2022). Similarly to the synthetic dataset, we have randomly selected some samples and removed the audio segment that follows a randomly chosen gap or pause.
3.3 Other details
Splits and statistics. The default train/dev/test split of the synthetic data follows that of the MultiWOZ dataset (Budzianowski et al., 2018). For the real data, we partition the collected YouTube and Buckeye conversations into disjoint train/dev/test splits at the file level, so that no source conversation appears in more than one split. The statistics of the synthetic and real datasets are given in Table˜2 and Table˜3, respectively.
Human validation. To quantify the label noise introduced by the automatic diarization + rule, we conducted a human validation study on a subset of real data (three annotators, clips). Annotators reached agreement with the automatic labels and rated diarization quality at on average; full results are in Appendix˜C.
@@ 4 Method
For a smooth real-time conversation, one needs to run the end-turn detector repetitively at a rapid frequency. Thus, it is desirable to deploy an ETD model with an extremely small computational footprint. However, lightweight models—e.g., a recurrent model with a single GRU layer—tend to suffer from a low detection accuracy. On the other hand, heavier transformer-based models typically require a large computational cost despite being highly accurate in the end-turn detection (see Section˜6).
To enjoy the best of both worlds, the SpeculativeETD is designed as a collaborative inference framework that utilizes both a lightweight model (say, GRU), which runs on-device, and a heavier model (say, Wav2vec 2.0), which runs on the server. In a nutshell, the proposed SpeculativeETD operates in two stages (Figure˜3):
@@
The key advantages of this framework are twofold: (1) The trickier decision that requires a large computation (i.e., telling “Gap” vs. “Pause”) can happen only once per each consecutive segment of silence, not at every time frame. As we will see in Section˜6, this helps save the computation more than 10 with only a small degradation prediction accuracy. (2) The two-stage pipeline allows us to utilize on-device computation by placing the lightweight model on the edge. By doing so, we no longer need the communication between the edge device and the server to be done in a continuous manner; instead, we can communicate once every silence.
@@
During real-time operation, the on-device model processes each new chunk in constant time. The server-side Wav2vec 2.0 is triggered once the on-device GRU predicts non-SU (silence) for at least , i.e., two consecutive chunks. Only the accumulated audio from the start of that silence segment is sent to the server. The server model then classifies the segment as Gap or Pause, determining whether an end-turn event has occurred or if streaming should continue. The trigger follows the standard turn-taking threshold (Stivers et al., 2009; Heldner and Edlund, 2010).
@@ 5 Experimental Setup
We now empirically validate the usefulness of the constructed OpenETD dataset (Section˜3) and the effectiveness of the proposed method, SpeculativeETD (Section˜4).
Task. We consider two tasks for the OpenETD:
@@
VAP (Ekstedt and Skantze, 2022): A popular open-source pretrained turn-taking model, which can also be employed for end-turn detection tasks. We use the pretrained encoder as frozen and train the predictor module of the model from scratch using the OpenETD dataset. We have used the “comparative” predictor head, which fits the ETD task.
GRU (Chung et al., 2014): We construct a lightweight GRU model with about parameters, consisting of a Conv2D frontend ( layers) and a single GRU layer on top of a -dim log–mel input computed on chunks (full input specification in Appendix˜B), and train it from scratch on the OpenETD dataset.
Wav2vec 2.0 (Baevski et al., 2020): A relatively large transformer-based model with 94M parameters, which achieves high performance in many speech processing tasks. We fine-tune the full model using the OpenETD dataset.
@@
In all experiments, we train the models on a mixture of the training splits of the synthetic and real data in the OpenETD dataset; the synthetic split provides controllable pause/gap patterns, while the real split grounds the models in natural conversational speech. For evaluation, we use the held-out test splits of both the synthetic and real datasets, with no overlap with the training or development sets.
@@
We have used the AdamW optimizer (Loshchilov and Hutter, 2019) for all tasks and models and trained for 10 epochs. The learning rate has been selected with a random search over the interval , using the validation split of the synthetic OpenETD dataset. The weight decay values are also randomly searched in the range . The batch size has been tuned over for GRU, for VAP, and for Wav2vec 2.0. The searched hyperparameters are provided in Appendix˜H.